OpenStudy (pottersheep):

4x^2/3 + 5x^1/3 - 6 = 0 What is the value of x?

OpenStudy (pottersheep):

I don't understand why there is only 1 answer....

OpenStudy (jamesj):

Let u = x^1/3 Then 4u^2 + 5u - 6 = 0, an equation that you can solve.

OpenStudy (pottersheep):

apparently the only aswer is 27/64 , but I get -8 as an answer as well

OpenStudy (pottersheep):

Thanks JamesJ I did that too, I got u = 3/4 and -2

myininaya (myininaya):

let u=x^(1/3) then u^2=x^(2/3) so we have 4u^2+5u-6=0 to factor we need to find two factors of 4(-6) that have product 4(-6) and sum 5 4(-6)=-12(2)=-3(8) 5=-3+8 so 5u=-3u+8u replace 5u with -3u+8u in 4u^2+5u-6=0 so we have 4u^2-3u+8u-6=0 now factor by grouping u(4u-3)+2(4u-3)=0 (4u-3)(u+2)=0 => u=3/4 or u=-2 but remember u=x^(1/3) so we have x^(1/3)=3/4 or x^(1/3)=-2 raise both sides to the third power x=(3/4)^3 or x=(-2)^(3) x=27/64 or x=-8 now we can check both of these by pluggin' them into the original equation so checking x=-8 4(-8)^(2/3)+5(-8)^(1/3)-6 = 4(-2)^2+5(-2)-6 = 4(4)-10-6 = 16-10-6 = 6-6 = 0 so both sides are 0! YAY!!! x=-8 works! now let's check x=27/64 4(27/64)^(2/3)+5(27/64)^(1/3)-6 = 4(3/4)^2+5(3/4)-6 = 4(9/16)+15/4-6 = 9/4+15/4-6 = 24/4-6 = 6-6 =0 both sides are zero that means x=27/64 works too! we have two solutions

OpenStudy (pottersheep):

THanks soooooooooo much =)) I guess the answer sheet is wrong. I became a fan too :))

myininaya (myininaya):

ok lol thanks

myininaya (myininaya):

does it ask for one solution?

myininaya (myininaya):

or does it tell you to list all solutions?